# Indoor Light and Heat Estimation from a Single Panorama

### v1

@@ 4.2 Sensitivity Analysis
The transient heat simulation (Equ. 3) accounts for three types of heat transfer: conduction, radiation, and convection. To evaluate the impact of each component on heat estimation, parametric heat simulations were performed with various combinations of these components, and corresponding heat maps for each condition are visualized, as shown in Fig. 11. When conduction, radiation, and convection are all included, the resulting heat map is considered the baseline for computing the error map. When there is no heat transfer (such as conduction, radiation, and convection), the temperature remains at its highest, as no heat is emitted from the object’s surface. Adding different types of heat transfer results in varying reductions in surface temperature. Each heat map is compared with the baseline image. In the error map, the first row shows a matrix with zero values, which indicates no temperature difference. Due to the low material conductivity in this simulation, the third row exhibits minimal temperature differences. Other combinations of heat transfer mechanisms increase the temperature by varying degrees, highlighting their individual and collective contributions to the heat distribution.
Various physical parameters in the heat simulation are analyzed for their impacts on indoor surface temperature. For a single panorama, we developed a graphic user interface (GUI) that allows users to click on the input image and outline circular target areas within indoor scenes. These selected target patches are used to compute the average temperature over their coverage, as shown in Fig. 12. We selected standard materials and parameters to generate the corresponding indoor heat maps. The default heat simulation time is 600, the surface thickness is 1, and the indoor ambient temperature is 21.11.
For different surface materials (Fig. 13), the surface temperatures and temperature differences (using the temperatures under plaster dense as baseline values) are visualized. Copper results in the lowest temperature on indoor surfaces, while plaster-dense, polystyrene, PVC (polyvinyl chloride), and simulated sheep wool result in higher temperatures than other materials and display similar temperature values across indoor surfaces.
Each material has its unique density, conductivity, and specific heat capacity. In this step, we set the default material to plaster-dense, adjust a single material property, and analyze its impact on surface temperature. Temperatures under different material densities and temperature differences (using the temperatures under correct density (1300 ) as baseline values) are visualized (Fig. 14). We observe that, as material density () increases, surface temperature decreases in general. While some temperatures rise on ceilings under 3300 and 5300 , with increments in material density, the temperatures at 1200, 1250, 1300, 1350, 1400, and 1500 are identical. When the material density reaches 7300 , it decreases the surface temperature by approximately between - 0.27 and 0.36 , with a corresponding percentage error between -1.00 and -1.60 %.
When changing material conductivity and specific heat capacity values, the surface temperature on indoor surfaces did not show differences (Fig. 15). This suggests that, compared to material density, conductivity and heat capacity have minor impacts on surface temperature when using plaster-dense as the material and our default simulation settings. While additional simulations could be conducted to further investigate this pattern, this section primarily focuses on the standard building materials and changes their thermal properties in the transient heat simulation.
Surface thickness is crucial in heat simulation, as thicker materials contain more mass, which alters the heat capacity. Following the default simulation settings, surface thickness is adjusted to examine its impact on surface temperatures. Temperatures under different surface thicknesses and their respective differences (using the temperatures for a 1 surface thickness as baseline values) are visualized in Fig. 16. As the surface thickness increases from 1 to 150 , the temperature continuously decreases. For example, for the surface temperature on the ceiling, when the thickness increased from 1 to 5 , the temperature dropped from 22.2 to below 21.6. However, when the thickness increased from 50 to 150 , the temperature change was minimal. Notably, for Wall 1, there was no temperature difference observed between thicknesses of 100 and 150 . Overall, when the indoor ambient temperature is 21.11, using thicker material results in a lower indoor temperature, bringing it closer to the ambient temperature in this setting.
When there is heat transfer between indoor and outdoor surfaces, Equ. 4 is used to estimate indoor heat maps at different outdoor temperatures. Fig. 17 shows the results with and without indoor-outdoor heat transfer at various outdoor temperatures. The image without indoor-outdoor heat transfer serves as the baseline for computing temperature differences. When the indoor ambient temperature is 21.11 , an outdoor temperature of 11.85 results in lower indoor surface temperatures. As the outdoor temperature rises above the indoor temperature, it leads to higher indoor surface temperatures, with greater heat transfer occurring from the outdoor environment to the indoor space. While the temperature difference is minor, using a temperature map ranging between -0.2 and 0.2 highlights the importance of considering indoor-outdoor temperature exchange for accurate indoor heat estimation in a real-world scenario.
4.3 Indoor Thermal Imaging
To evaluate the results of our heat simulation, we captured the real-world scenes using a Ricoh Theta Z1 camera alongside a thermal camera. The heat simulation enables visualization of indoor heat maps in the equirectangular representation, while the thermal camera provides only a 2D perspective view. Therefore, we cropped the target region from the rendered heat map and converted it into a 2D perspective with the same field of view (FOV) as the thermal camera. We then evaluated the rendered images by comparing them with the thermal images captured by the thermal camera.
As shown in Fig. 18, a Ricoh Theta Z1 camera is positioned side by side with the FLIR thermal camera at the same height, with a Macbeth color checker and a gray matte board placed within the field of view. The real-time indoor temperature is recorded on the Govee Hygrometer Thermometer (H5075), and the outdoor temperature is obtained from GPS data in the local region.
We set up the FLIR thermal camera in TLinear mode (Boson, image resolution: 640 by 512 pixels) when capturing the thermal images. However, TLinear assumes the default emissivity, atmosphere temperature, and background temperature for the scene. The thermal image captured by TLinear requires a post-process (Veitch-Michaelis (2024)) to display the correct thermal image based on the actual emissivity of the material.
The counts () from the digital converter (ADC) can be expressed as:
Where is the thermal image captured under TLinear mode, , , , and are FLIR’s constant calibration coefficients.
The flux () at the detector can be expressed as:
Where represents the default emissivity value, and is the default background temperature (a constant value).
The actual actual emissivity () is used to compute the :
The thermal image, displayed with the actual material emissivity, can be computed as:
@@ 4.4 Validation and Comparison
Using our proposed heat estimation method, a single indoor panorama is used to generate a panoramic heat map from the same camera position (Fig. 19). The average temperature values are calculated on the planar indoor surfaces within the white circles, and the heat estimation results are compared with the thermal images captured by the thermal camera. In Scene 1, the temperature difference at selected areas ranged from 2.00 to 2.50 . In Scene 2, the temperature difference was approximately 4.00 to 4.50 . In Scene 3, the wall plane in the upper left corner showed a temperature difference of about 1.00 , while the wall plane on the right and the floor exhibited larger temperature errors, close to 2.00 . It is also important to note that the thermal images from the thermal camera are noisy, which can lower the average temperature in the target region.
An accurate 3D floor layout is essential for precise heat estimation. We focus on the error analysis of the heat simulation in different floor layouts (Fig. 20). Layout 1 is considered as the correct floor layout, while Layouts 2 and 3 feature slight shifts in the floor boundaries. The heat maps for Layouts 2 and 3 reveal variations in indoor temperature values due to inaccuracies in the room geometry. In the Error Map, the heat maps of Layouts 2 and 3 are compared with the ground-truth (GT) heat map of Layout 1. Layout 2 shows lower temperature estimates at the upper and lower corners of the wall, while Layout 3 exhibits lower temperatures near the wall adjacent to the left window. Although the temperature differences are minor, an accurate floor layout is crucial to generate reliable indoor heat maps.
@@ 5 Conclusion
In this work, we present an application that uses paired indoor and outdoor RGB panoramas as input to estimate indoor lighting and heat properties. Built on image-based rendering techniques, our inverse rendering pipeline not only provides high-quality indoor renderings in a photorealistic manner but also enables us to estimate light levels for indoor task lighting analysis. Since the panorama provides an omnidirectional view of the scene, the task lighting tool allows users to identify suitable work areas. Besides, heat estimation is integrated with lighting analysis. The heat generated by natural light on indoor surfaces is estimated, whereas transient heat transport enables the analysis of indoor heat distribution over time. Our method provides an efficient solution for indoor lighting and thermal analysis, supporting virtual indoor staging and contributing to high-quality scene design, lighting research, and thermal imaging techniques for real-world scenarios.
@@ 6.1 Limitations
This paper presents the methodology for a joint estimation of light and heat for a single panorama. However, several limitations are identified in this work. The heat estimation assumes that all indoor heat originates from natural light (the outdoor RGB panorama) and takes sunlight as the sole heat source. Since mechanical systems, such as air conditioning or heating operations, are unknown in a single panorama, their power and functionality remain unknown. Consequently, the heat estimation reflects only the temperature effects caused by outdoor light. The outdoor scene is captured as a 360∘ environment map. This map serves as an omnidirectional illumination source from an infinite distance. However, accurately modeling detailed outdoor 3D objects can be challenging when obstructions are present. For example, under clear sky conditions, sunlight can be blocked by trees and leaves, resulting in light leakage as beams pass through the foliage. These subtle lighting effects cannot be recovered effectively, as the scene is flattened onto the environment map. The estimation of irradiance (illuminance) is based on the assumption that all objects have Lambertian surfaces. Materials such as glass or metal have not been included in this work.
@@ 6.2 Future Work
7 Disclosure Statement
The authors declare no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.

### v2

@@ 3.4 Sensitivity Analysis
The transient heat simulation (Eq. 3) accounts for three types of heat transfer: conduction, radiation, and convection. To evaluate the impact of each component on heat estimation, parametric heat simulations were performed with various combinations of these components, and the corresponding heat maps for each condition are visualized, as shown in Fig. 8. When conduction, radiation, and convection are all included, the resulting heat map is considered the baseline for computing the error map. When there is no heat transfer (such as conduction, radiation, and convection), the temperature remains at its highest, as no heat is emitted from the object’s surface. Adding different types of heat transfer results in varying reductions in surface temperature. Each heat map is compared with the baseline image (first row, left column). In the error maps, the baseline image displays a matrix of zero values, indicating the absence of temperature differences. Owing to the low material conductivity in the heat simulation, the image in the third row, left column exhibits minimal temperature variation. Other combinations of heat transfer mechanisms increase the temperature to varying degrees, highlighting their individual and collective contributions to the overall heat distribution.
@@ 3.5 Validation and Comparison
To evaluate the results of our heat simulation, we captured real-world scenes using a Ricoh Theta Z1 camera alongside a thermal camera. The heat simulation enables visualization of indoor heat maps in the equirectangular representation, while the thermal camera provides only a 2D perspective view. Therefore, we approximately cropped the target region from the rendered heat map and converted it into a 2D perspective with the same field of view (FOV) as the thermal camera. We then evaluated the rendered images by comparing them with the thermal images captured by the thermal camera.
When capturing visible-light and thermal images, a Ricoh Theta Z1 camera is positioned side by side with the FLIR thermal camera at the same height, with a Macbeth color checker and a gray matte board placed within the field of view. The real-time indoor temperature is recorded on the Govee Hygrometer Thermometer (H5075), and the outdoor temperature is obtained from GPS data and local time.
We set up the FLIR thermal camera in TLinear mode (Boson, image resolution: 640 by 512 pixels) when capturing the thermal images. However, TLinear mode assumes the default emissivity, atmosphere temperature, and background temperature for the scene. The thermal image captured by TLinear requires a post-process [53] to display the correct thermal image based on the actual emissivity of the materials.
Using our proposed heat estimation method, a single indoor panorama is used to generate a panoramic heat map from the same camera position (Fig. 9). The average temperature values are calculated on the planar indoor surfaces within the white circles, and the heat estimation results are compared with the thermal images captured by the thermal camera. In Scene 1, the temperature difference in selected areas ranged from around 2.00 to 2.50 . In Scene 2, the temperature difference was approximately 4.00 to 4.50 . In Scene 3, the wall plane in the upper left corner showed a temperature difference of about 1.00 , while the wall plane on the right and the floor exhibited larger temperature errors, close to 2.00 . The simulation discrepancy arises from several factors, such as complex heat sources (air conditioning, electronic appliances), uncertainties in material properties, and the simplified Lambertian surface assumption used to model the real-world scene. It is also important to note that the thermal images from the thermal camera are usually noisy, which can lower the average temperature in the target region.
@@ 4 Conclusion
In this work, we present an application that uses paired indoor and outdoor RGB panoramas as input to estimate indoor lighting and heat properties. Built on image-based rendering techniques, our inverse rendering pipeline not only provides high-quality indoor renderings in a photorealistic manner but also enables us to estimate light levels for indoor task lighting analysis. Since the panorama provides an omnidirectional view of the scene, our application allows users to identify suitable work areas. In addition, heat estimation is integrated with light estimation. The heat generated by natural light on indoor surfaces is estimated, whereas transient heat transport enables the analysis of indoor heat distribution over time. Our method provides an efficient solution for indoor lighting and thermal analysis, supporting virtual indoor staging and contributing to high-quality scene visualization, lighting research, and thermal imaging techniques for real-world scenarios.
@@ 5.1 Limitations
This paper presents the methodology for a joint estimation of light and heat for a single visible-light panorama. However, several limitations are identified in this work. The heat estimation assumes that all indoor heat originates from natural light (the outdoor RGB panorama) and takes sunlight as the sole heat source. Since mechanical systems, such as air conditioning or heating operations, are unknown in a single visible-light panorama, their power and functionality remain unknown. Consequently, the heat estimation reflects only the temperature effects caused by outdoor light. The outdoor scene is captured as a 360∘ environment map. This map serves as an omnidirectional illumination source from an infinite distance. However, accurately modeling detailed outdoor 3D objects can be challenging when obstructions are present. For example, under clear sky conditions, sunlight can be blocked by trees and leaves, resulting in light leakage as beams pass through the foliage. These subtle lighting effects cannot be recovered effectively, as the scene is flattened onto the environment map. The estimation of irradiance (illuminance) is based on the assumption that all objects have Lambertian surfaces. Materials such as glass or metal have not been included in this work.
@@ 5.2 Future Work
Acknowledgement
This work was partially supported by a gift from Zillow Group, USA, and NSF grant IIS2107236.